# Elevated UV luminosity density at Cosmic Dawn explained by non-evolving, weakly-mass dependent star formation efficiency

### v1

@@ Abstract
pubyear: 2024pagerange: Elevated UV luminosity density at Cosmic Dawn explained by non-evolving, weakly-mass dependent star formation efficiencyD
@@ 1 Introduction
The advent of JWST has transformed our understanding of high-redshift galaxy evolution thanks to its excellent sensitivity, resolution, and wavelength coverage. Among many other findings, JWST revealed a higher-than-expected density of UV bright galaxies at high $z$ (Naidu et al., 2022; Finkelstein et al., 2022; Labbé et al., 2023; Finkelstein et al., 2023; Casey et al., 2024), confirmed spectroscopically the presence of galaxies up to $z\sim{}14$ (Curtis-Lake et al., 2023; Robertson et al., 2023b; Harikane et al., 2024; Castellano et al., 2024; Carniani et al., 2024), and uncovered the population of galaxies most likely responsible for reionization (Atek et al., 2024). Furthermore, JWST-based studies showed an elevated UV luminosity density $\rho_{\rm UV}$ at $z>10$ compared with both previous observational estimates and modeling predictions (Donnan et al., 2023a; Donnan et al., 2023b; Harikane et al., 2024; Harikane et al., 2023; Finkelstein et al., 2023; Bouwens et al., 2023b) suggesting that an important component of galaxy theory is either missing or not sufficiently understood. Solving this conundrum is critical, not only to gain a better understanding of the physical processes driving galaxy evolution at high $z$, but also to constrain the impact of galaxies on cosmic reionization.
@@
First, models with a non-evolving SFE – halo mass relation match the observed UV luminosity density at $z\sim{}6-10$ but underpredict $\rho_{\rm UV}$ at $z>10$ (e.g., Mason et al. 2015; Tacchella et al. 2018; Harikane et al. 2022). This argument is further strengthened by many empirical or semi-analytical models appearing to favor significant evolution to match observations (Sun & Furlanetto, 2016; Behroozi et al., 2019; Sipple & Lidz, 2024; Chakraborty & Choudhury, 2024). Cosmological hydrodynamical simulations targeting galaxies at $z>10$ have not fully resolved this issue as they yield significantly varying predictions for the UV luminosity density (e.g., Wilkins et al. 2022; Kannan et al. 2023; Ceverino et al. 2024), perhaps indicating gaps in our understanding of the physics during the Cosmic Dawn. Intriguingly, recent simulations analyzing $z=8-12$ suggest that this tension could be resolved with standard astrophysical modeling (Sun et al., 2023c).
Secondly, the observation of massive, luminous galaxies suggests that at least some galaxies are efficient in converting accreted baryonic matter into stars (Labbé et al., 2023; Boylan-Kolchin, 2023; Casey et al., 2024), potentially pointing to a failure of feedback processes limiting star formation (Bassini et al., 2023; Dekel et al., 2023; Li et al., 2023; Mirocha & Furlanetto, 2023). We note, however, that some observational studies suggest a mild to absent evolution in the SFE of galaxies at high redshift (Stefanon et al., 2021; Harikane et al., 2022). Moreover, the presence of UV bright galaxies may also be explained in other ways, e.g., via a highly stochastic star formation activity (Sun et al., 2023a; Pallottini & Ferrara, 2023; Sun et al., 2023c; Shen et al., 2023a; Kravtsov & Belokurov, 2024), low dust attenuation (Ferrara et al., 2023), or a top-heavy IMF that increases the UV luminosity per SFR (Harikane et al., 2023).
@@ 2.2 Post-processing
Figure 1: Rest-frame ultraviolet luminosity functions (UV-LFs) of galaxies in FIREboxHR at $z\sim{}6-15$ (colored circles and lines) and comparison with observational data (black symbols and thick lines; thin black lines show extrapolations of the observed UV-LFs to faint magnitudes): McLure et al. 2013 (L13), Harikane et al. 2022 (H22), Naidu et al. 2022 (N22), Bouwens et al. 2022 (B22), Donnan et al. 2023a (D23), Harikane et al. 2023 (H23), Pérez-González et al. 2023 (P23), Bouwens et al. 2023b (B23), Atek et al. 2024 (A24), Casey et al. 2024 (C24), Finkelstein et al. 2024 (F24), Robertson et al. 2023a (R24), and Donnan et al. 2024 (D24). The UV luminosities are calculated with the dust-radiative transfer code SKIRT for the more massive galaxies (filled circles) and in an optically thin approximation for the remaining, largely dust-free, galaxies (empty circles). Vertical error bars of the UV-LFs in FIREboxHR are obtained via bootstrapping, while horizontal error bars indicate bin-widths. Thinner colored lines indicate the UV-LFs when galaxies with stellar masses below $10^{5}\,M_{\odot}$ ($\sim{}12-15$ star particles) or below $3\times{}10^{5}\,M_{\odot}$ ($\sim{}35-50$ star particles) are excluded. FIREboxHR reproduces the observed UV LFs and their evolution over $z\sim{}6-14$.
@@ 3.1 UV luminosity function and luminosity density in FIREboxHR
According to Fig. 1, the UV LFs predicted by FIREboxHR are in good agreement with observational estimates from JWST and HST at $z\sim{}6-14$ (McLure et al., 2013; Harikane et al., 2022; Naidu et al., 2022; Bouwens et al., 2022; Donnan et al., 2023a; Harikane et al., 2023; Pérez-González et al., 2023; Bouwens et al., 2023b; Robertson et al., 2023a; Atek et al., 2024; Finkelstein et al., 2024; Donnan et al., 2024). In particular, the LFs follow closely the observed double-power-law or Schechter function parametrization over a broad range in UV luminosity and with a normalization that decreases towards higher redshifts.
@@
The role of dust attenuation is limited to the bright end of our UV LFs. For instance, the number density of galaxies at $z=6.3-10$ with a UV magnitude of $-18.5$ increases by about $0.1-0.3$ dex if we consider intrinsic, i.e., not dust-attenuated UV luminosities, while the LF at UV magnitudes fainter than -17.5 is unchanged. The presence or absence of dust also has no impact on our UV LFs at $z>10$ given the dearth of massive, bright galaxies in the simulation at those redshifts.
In this work we calculate the cosmic UV luminosity density by adding the UV luminosities of all galaxies brighter than a given UV magnitude limit $M_{\rm UV,lim}$ and then dividing the total luminosity by the simulation volume. This direct approach is feasible since FIREboxHR reproduces the number and luminosities of galaxies dominating the cosmic UV luminosity density. Specifically, as evidenced by Fig. 2, half the dust-attenuated UV luminosity density in galaxies brighter than $M_{\rm UV,lim}=-17$, denoted as $M_{\rm UV,50\%}$, arises from the subset of galaxies fainter than -18.4 at $z=6.3$ and fainter than about -18 at $z=10$. These UV luminosities correspond to galaxies with typical stellar masses of $M_{\rm star}\sim{}10^{7.5}-10^{8}\,M_{\odot}$ that reside in halos with $M_{\rm halo}\sim{}10^{10}-10^{10.5}\,M_{\odot}$, see Fig. 13 in the appendix. These stellar masses are in line with observations, e.g., Endsley et al. (2023a) find that galaxies with $M_{\rm UV}\sim{}-18.4$ have typical stellar masses of $M_{\rm star}\sim{}10^{7.5}-10^{8}\,M_{\odot}$ at $z\sim{}6-7$.
@@
Figure 3: Evolution of the dust-attenuated ultraviolet (UV) luminosity density in FIREboxHR at $z\sim{}6-15$ for different values of the UV magnitude limit. Uncertainties (16-84%) in the UV luminosity density are estimated via bootstrapping. The simulation data is compared both with observational data (Bouwens et al., 2022; Donnan et al., 2023a; Donnan et al., 2023b; Harikane et al., 2023) as well as with theoretical predictions based on the empirical models by Tacchella et al. (2018) and Harikane et al. (2022), the Cosmic re-ionization on computers (CROC) simulation suite (Gnedin, 2014), FIRE-2 zoom-in simulation suites by Ma et al. (2019) and Sun et al. (2023c), the SPHINX20 simulation (Rosdahl et al., 2022), the THESAN-1 simulation (Kannan et al., 2022), the Millenium-TNG simulation (Kannan et al., 2023), and the FirstLight zoom-in suite (Ceverino et al., 2024). A UV magnitude limit of -17 is adopted for all literature data. The treatment of dust attenuation in the various models is discussed in the text. The UV luminosity density predicted by FIREboxHR agrees well with observational data, in particular those based on recent JWST observations, and it evolves more gradually at high redshift than predicted by many previous models.
Fig. 3 shows the cosmic UV luminosity density in FIREboxHR calculated as described above for different choices of $M_{\rm UV,lim}$. We compare the simulation predictions to observational data with JWST and HST for the common choice of $M_{\rm UV,lim}=-17$ (Bouwens et al., 2015; Donnan et al., 2023a; Donnan et al., 2023b; Harikane et al., 2023) finding excellent agreement. The simulation results are also consistent with recent observations by Finkelstein et al. (2024) and Donnan et al. (2024). The apparent upturn at $z=15$ in FIREboxHR is a result of low number statistics combined with a burst of star formation a few Myr earlier boosting UV luminosity. The figure also shows predictions for $M_{\rm UV,lim}>-17$. At $z\lesssim{}10$, the cosmic UV luminosity density is not very sensitive to the choice of the limiting magnitude provided the latter is chosen to be fainter than about -14. For instance, the UV luminosity density at $z=8$ increases by 0.06 dex (0.30 dex) if $M_{\rm UV,lim}$ is increased from -14 to -12 (from -17 to -14). We attribute the insensitivity of the cosmic UV luminosity to the UV magnitude limit at $z\lesssim{}10$ to the shallow slope (turn-over) of the faint end of the UV LFs shown in Fig. 1. In contrast, at $z\gtrsim{}11$, i.e., before the onset of reionization, the cosmic UV luminosity depends much more noticeably on the chosen UV magnitude limit. For example, at $z=13.5$, the cosmic UV luminosity density changes by 0.29 dex if $M_{\rm UV,lim}$ is raised from -14 to -12.
@@
Fig. 3 also includes the UV luminosity density predicted by constant efficiency models (Tacchella et al., 2018; Harikane et al., 2022), the Cosmic reionization on computers (CROC) simulation suite (Gnedin, 2014), suites of zoom-in FIRE simulations by Ma et al. (2019) and Sun et al. (2023c), the SPHINX20 simulation (Rosdahl et al., 2022), the THESAN-1 simulation (Kannan et al., 2022), the Millenium-TNG simulation (Kannan et al., 2023), and the FirstLight zoom-in suite (Ceverino et al., 2024), all for $M_{\rm UV,lim}=-17$. The FLARES simulation (Wilkins et al., 2022) reports results only for galaxies brighter than about -18 and is not included in this comparison.
@@
We convert the estimates of the cosmic SFR density by Tacchella et al. (2018) and Harikane et al. (2022) to a luminosity density with the help of the Madau & Dickinson (2014) conversion factor. To account for dust attenuation, we first calculate the typical UV magnitude of galaxies contribution to the cosmic luminosity density via eq. (1). We then follow Tacchella et al. (2018) to derive the expected $A_{\rm V}$ of such galaxies via the method proposed by Smit et al. (2012) and reduce the UV luminosity density accordingly. This dust correction primarily affects results for $z\leq{}7$. For the CROC suite, we numerically integrate the UV LF functions shown in their Fig. 1 to a magnitude limit of -17. These LFs account for dust attenuation in an approximate fashion based on the overall stellar metallicity each galaxy and a redshift dependent dust column density calibrated against observations (Bouwens et al., 2009). The luminosity densities for SPHINX and THESAN are obtained similarly by sampling the provided UV LFs and integrating them to a magnitude limit of -17. The LFs in THESAN account for dust attenuation in a similar way as Gnedin (2014) but with the redshift-dependent dust-to-metal ratio from Vogelsberger et al. (2020), while those from SPHINX are calculated with dust-radiative transfer. The UV luminosity density provided by Ma et al. (2019) accounts for dust attenuation by post-processing with SKIRT. The results of Sun et al. (2023c) are based on integrating their fitted UV LFs at $z=8-12$ over the UV magnitude range -26 and -17 with dust attenuation calculated as described in their article, i.e., by combining the $A_{\rm UV}-\beta_{\rm UV}$ relation by Fudamoto et al. (2020) with the $\beta_{\rm UV}-M_{\rm UV}$ relation by Cullen et al. (2023). Similarly, we obtain an estimate of the UV luminosity density in Millenium-TNG by integrating the provided Schechter functions fits of the dust-attenuated UV LFs over the UV magnitude range -26 and -17. Kannan et al. (2023) employ an empirical dust-attenuation model fitted to ALMA observations of galaxies at $z\sim{}4-7$ by Bouwens et al. (2016b). Finally, Ceverino et al. (2024) adopt an empirical dust correction based on the observed relation between UV continuum slope and UV magnitude at $z\sim{}6-9$ (Bouwens et al., 2014; Atek et al., 2023).
@@
While these models all predict a UV luminosity density in approximate agreement with both FIREboxHR and observational data at $z\sim{}6-9$, most of them increasingly underestimate the UV luminosity density towards higher redshift. Exceptions are the SPHINX20 simulation (Rosdahl et al., 2022), which matches the predictions of FIREboxHR at $z=10$ but does not provide estimates for higher redshifts, the FIRE-2 zoom-in suites by Ma et al. (2019) and Sun et al. (2023c), which match FIREboxHR out to $z=10$ and $z=12$, respectively, and the FirstLight suite (Ceverino et al., 2024) which underpredicts the observed UV luminosity density at $\sim{}10-11$ but is in approximate agreement at $z\sim{}12-13$.

### v2

# Elevated UV luminosity density at Cosmic Dawn explained by non-evolving, weakly mass-dependent star formation efficiency

@@ Abstract
pubyear: 2024pagerange: Elevated UV luminosity density at Cosmic Dawn explained by non-evolving, weakly mass-dependent star formation efficiencyE
@@ 1 Introduction
The advent of JWST has transformed our understanding of high-redshift galaxy evolution thanks to its excellent sensitivity, resolution, and wavelength coverage. Among many other findings, JWST revealed a higher-than-expected density of UV bright galaxies at high $z$ (Naidu et al., 2022; Finkelstein et al., 2022; Labbé et al., 2023; Finkelstein et al., 2023; Casey et al., 2024), confirmed spectroscopically the presence of galaxies up to $z\sim{}14$ (Curtis-Lake et al., 2023; Robertson et al., 2023; Harikane et al., 2024; Castellano et al., 2024; Carniani et al., 2024), and uncovered the population of galaxies most likely responsible for reionization (Atek et al., 2024). Furthermore, JWST-based studies showed an elevated UV luminosity density $\rho_{\rm UV}$ at $z>10$ compared with both previous observational estimates and modeling predictions (Donnan et al., 2023a; Donnan et al., 2023b; Harikane et al., 2024; Harikane et al., 2023; Finkelstein et al., 2023; Bouwens et al., 2023b; Chemerynska et al., 2024; Conselice et al., 2024) suggesting that an important component of galaxy theory is either missing or not sufficiently understood. Solving this conundrum is critical, not only to gain a better understanding of the physical processes driving galaxy evolution at high $z$, but also to constrain the impact of galaxies on cosmic reionization.
@@
First, models with a non-evolving SFE – halo mass relation match the observed UV luminosity density at $z\sim{}6-10$ but underpredict $\rho_{\rm UV}$ at $z>10$ (e.g., Mason et al. 2015; Tacchella et al. 2018; Harikane et al. 2022). This argument is further strengthened by many empirical or semi-analytical models appearing to favor significant evolution to match observations (Sun & Furlanetto, 2016; Behroozi et al., 2019; Sabti et al., 2022; Sipple & Lidz, 2024; Qin et al., 2023; Chakraborty & Choudhury, 2024; Wang et al., 2024). Cosmological hydrodynamical simulations targeting galaxies at $z>10$ have not fully resolved this issue as they yield significantly varying predictions for the UV luminosity density (e.g., Wilkins et al. 2022; Kannan et al. 2023; Ceverino et al. 2024), perhaps indicating gaps in our understanding of the physics during the Cosmic Dawn. Intriguingly, recent simulations analyzing $z=8-12$ suggest that this tension could be resolved with standard astrophysical modeling (Sun et al., 2023c).
Secondly, the observation of massive, luminous galaxies suggests that at least some galaxies are efficient in converting accreted baryonic matter into stars (Labbé et al., 2023; Boylan-Kolchin, 2023; Casey et al., 2024), potentially pointing to a failure of feedback processes limiting star formation (Bassini et al., 2023; Dekel et al., 2023; Li et al., 2024; Mirocha & Furlanetto, 2023). We note, however, that some observational studies suggest a mild to absent evolution in the SFE of galaxies at high redshift (Stefanon et al., 2021; Harikane et al., 2022). Moreover, the presence of UV bright galaxies may also be explained in other ways, e.g., via a highly stochastic star formation activity (Sun et al., 2023a; Pallottini & Ferrara, 2023; Sun et al., 2023c; Shen et al., 2023a; Kravtsov & Belokurov, 2024), low dust attenuation (Ferrara et al., 2023), AGN contamination (Hegde et al., 2024), or a top-heavy IMF that increases the UV luminosity per SFR (Harikane et al., 2023).
@@ 2.2 Post-processing
Figure 1: Rest-frame ultraviolet luminosity functions (UV-LFs) of galaxies in FIREboxHR at $z\sim{}6-15$ (colored circles, squares, and lines) and comparison with observational data (black symbols and thick lines; thin black lines show extrapolations of the observed UV-LFs to faint magnitudes): McLure et al. 2013 (L13), Harikane et al. 2022 (H22), Naidu et al. 2022 (N22), Bouwens et al. 2022 (B22), Bouwens et al. 2023b (B23), Donnan et al. 2023a (D23), Harikane et al. 2023 (H23), Pérez-González et al. 2023 (P23), Adams et al. 2024 (Ad24), Atek et al. 2024 (At24), Casey et al. 2024 (C24), Donnan et al. 2024 (D24), Finkelstein et al. 2024 (F24), and Robertson et al. 2024 (R24). The UV luminosities are calculated with the dust-radiative transfer code SKIRT for the more massive galaxies (filled circles) and in an optically thin approximation for the remaining, largely dust-free, galaxies (empty circles). Vertical error bars of the UV-LFs in FIREboxHR are obtained via bootstrapping, while horizontal error bars indicate bin-widths. Thinner colored lines indicate the UV-LFs when galaxies with stellar masses below $10^{5}\,M_{\odot}$ ($\sim{}12-15$ star particles) or below $3\times{}10^{5}\,M_{\odot}$ ($\sim{}35-50$ star particles) are excluded. FIREboxHR reproduces the observed UV LFs at moderate to low luminosities and their evolution over $z\sim{}6-14$.
@@ 3.1 UV luminosity function and luminosity density in FIREboxHR
According to Fig. 1, the UV LFs predicted by FIREboxHR are in good agreement with observational estimates from JWST and HST at $z\sim{}6-14$ (McLure et al., 2013; Harikane et al., 2022; Naidu et al., 2022; Bouwens et al., 2022; Donnan et al., 2023a; Harikane et al., 2023; Pérez-González et al., 2023; Bouwens et al., 2023b; Robertson et al., 2024; Adams et al., 2024; Atek et al., 2024; Finkelstein et al., 2024; Donnan et al., 2024). In particular, the LFs follow closely the observed double-power-law or Schechter function parametrization over a broad range in UV luminosity and with a normalization that decreases towards higher redshifts.
@@
The role of dust attenuation is limited to the bright end of our UV LFs. For instance, the number density of galaxies at $z=6.3-10$ with a UV magnitude of $-18.5$ increases by about $0.1-0.3$ dex if we consider intrinsic, i.e., not dust-attenuated UV luminosities, while the LF at UV magnitudes fainter than -17.5 is unchanged. The presence or absence of dust also has no impact on our UV LFs at $z>10$ given the dearth of massive, bright galaxies in the simulation at those redshifts. The limited dust attenuation in FIREboxHR galaxies reflects their low stellar masses and, consequently, low gas-phase metallicities. For a detailed analysis of the mass – metallicity relation in high-redshift galaxies with FIRE-2 physics, see Marszewski et al. (2024).
In this work we calculate the cosmic UV luminosity density by adding the UV luminosities of all galaxies brighter than a given UV magnitude limit $M_{\rm UV,lim}$ and then dividing the total luminosity by the simulation volume. This direct approach is feasible since FIREboxHR reproduces the number and luminosities of galaxies dominating the cosmic UV luminosity density. Specifically, as evidenced by Fig. 2, half the dust-attenuated UV luminosity density in galaxies brighter than $M_{\rm UV,lim}=-17$, denoted as $M_{\rm UV,50\%}$, arises from the subset of galaxies fainter than -18.4 at $z=6.3$ and fainter than about -18 at $z=10$. These UV luminosities correspond to galaxies with typical stellar masses of $M_{\rm star}\sim{}10^{7.5}-10^{8}\,M_{\odot}$ that reside in halos with $M_{\rm halo}\sim{}10^{10}-10^{10.5}\,M_{\odot}$, see Fig. 13 in the appendix. These stellar masses are in line with observations, e.g., Endsley et al. (2024) find that galaxies with $M_{\rm UV}\sim{}-18.4$ have typical stellar masses of $M_{\rm star}\sim{}10^{7.5}-10^{8}\,M_{\odot}$ at $z\sim{}6-7$.
@@
Figure 3: Evolution of the dust-attenuated ultraviolet (UV) luminosity density in FIREboxHR at $z\sim{}6-15$ for different values of the UV magnitude limit. Uncertainties (16-84%) in the UV luminosity density are estimated via bootstrapping. The simulation data is compared with observational data by Bouwens et al. (2022), Donnan et al. (2023a); Donnan et al. (2023b), Harikane et al. (2023), Adams et al. (2024), Finkelstein et al. (2024) and Donnan et al. (2024), as well as with theoretical predictions based on the empirical models by Tacchella et al. (2018) and Harikane et al. (2022), the Cosmic re-ionization on computers (CROC) simulation suite (Gnedin, 2014), FIRE-2 zoom-in simulation suites by Ma et al. (2019) and Sun et al. (2023c), the SPHINX20 simulation (Rosdahl et al., 2022), the THESAN-1 simulation (Kannan et al., 2022), the Millennium-TNG simulation (Kannan et al., 2023), and the FirstLight zoom-in suite (Ceverino et al., 2024). A UV magnitude limit of -17 is adopted for all literature data. Small horizontal shifts ($\Delta{}z<0.05$) are applied to some data points to increase readability. The treatment of dust attenuation in the various models is discussed in the text. The UV luminosity density predicted by FIREboxHR agrees well with observational data, in particular those based on recent JWST observations, and it evolves more gradually at high redshift than predicted by many previous models.
Fig. 3 shows the cosmic UV luminosity density in FIREboxHR calculated as described above for different choices of $M_{\rm UV,lim}$. We compare the simulation predictions to observational data from JWST and HST for the common choice of $M_{\rm UV,lim}=-17$ (Bouwens et al., 2015; Donnan et al., 2023a; Donnan et al., 2023b; Harikane et al., 2023; Adams et al., 2024; Finkelstein et al., 2024; Donnan et al., 2024) finding overall good agreement. At $z\sim{}10-11$, the predicted UV luminosity density is in line with Donnan et al. (2023a); Donnan et al. (2023b); Donnan et al. (2024) and Finkelstein et al. (2024) but higher ($\sim{}0.3$ dex) than the results reported by Adams et al. (2024). The apparent upturn at $z=15$ in FIREboxHR is a result of low number statistics combined with a burst of star formation a few Myr earlier boosting UV luminosity.
The figure also shows predictions for UV magnitude limits fainter than -17. At $z\lesssim{}10$, the cosmic UV luminosity density is not very sensitive to the choice of the limiting magnitude provided the latter is chosen to be fainter than about -14. For instance, the UV luminosity density at $z=8$ increases by 0.06 dex (0.30 dex) if $M_{\rm UV,lim}$ is increased from -14 to -12 (from -17 to -14). We attribute the insensitivity of the cosmic UV luminosity to the UV magnitude limit at $z\lesssim{}10$ to the shallow slope (turn-over) of the faint end of the UV LFs shown in Fig. 1. In contrast, at $z\gtrsim{}11$, i.e., before the onset of reionization, the cosmic UV luminosity depends much more noticeably on the chosen UV magnitude limit. For example, at $z=13.5$, the cosmic UV luminosity density changes by 0.29 dex if $M_{\rm UV,lim}$ is raised from -14 to -12.
@@
Fig. 3 also includes the UV luminosity density predicted by constant (i.e., non-evolving) efficiency models (Tacchella et al., 2018; Harikane et al., 2022), the Cosmic reionization on computers (CROC) simulation suite (Gnedin, 2014), suites of zoom-in FIRE simulations by Ma et al. (2019) and Sun et al. (2023c), the SPHINX20 simulation (Rosdahl et al., 2022), the THESAN-1 simulation (Kannan et al., 2022), the Millenium-TNG simulation (Kannan et al., 2023), and the FirstLight zoom-in suite (Ceverino et al., 2024), all for $M_{\rm UV,lim}=-17$. The FLARES simulation (Lovell et al., 2020; Wilkins et al., 2022) reports results only for galaxies brighter than about -18 and is not included in this comparison.
@@
We convert the estimates of the cosmic SFR density by Tacchella et al. (2018) and Harikane et al. (2022) to a luminosity density with the help of the Madau & Dickinson (2014) conversion factor. To account for dust attenuation, we first calculate the typical UV magnitude of galaxies contribution to the cosmic luminosity density via Eq. (1). We then follow Tacchella et al. (2018) to derive the expected $A_{\rm V}$ of such galaxies via the method proposed by Smit et al. (2012) and reduce the UV luminosity density accordingly. This dust correction primarily affects results for $z\leq{}7$. For the CROC suite, we numerically integrate the UV LF functions shown in their Fig. 1 to a magnitude limit of -17. These LFs account for dust attenuation in an approximate fashion based on the overall stellar metallicity each galaxy and a redshift dependent dust column density calibrated against observations (Bouwens et al., 2009). The luminosity densities for SPHINX and THESAN are obtained similarly by sampling the provided UV LFs and integrating them to a magnitude limit of -17. The LFs in THESAN account for dust attenuation in a similar way as Gnedin (2014) but with the redshift-dependent dust-to-metal ratio from Vogelsberger et al. (2020), while those from SPHINX are calculated with dust-radiative transfer. The UV luminosity density provided by Ma et al. (2019) accounts for dust attenuation by post-processing with SKIRT. The results of Sun et al. (2023c) are based on integrating their fitted UV LFs at $z=8-12$ over the UV magnitude range -26 and -17 with dust attenuation calculated as described in their article, i.e., by combining the $A_{\rm UV}-\beta_{\rm UV}$ relation by Fudamoto et al. (2020) with the $\beta_{\rm UV}-M_{\rm UV}$ relation by Cullen et al. (2023). Similarly, we obtain an estimate of the UV luminosity density in Millennium-TNG by integrating the provided Schechter functions fits of the dust-attenuated UV LFs over the UV magnitude range -26 and -17. Kannan et al. (2023) employ an empirical dust-attenuation model fitted to ALMA observations of galaxies at $z\sim{}4-7$ by Bouwens et al. (2016b). Finally, Ceverino et al. (2024) adopt an empirical dust correction based on the observed relation between UV continuum slope and UV magnitude at $z\sim{}6-9$ (Bouwens et al., 2014; Atek et al., 2023).
@@
While these models all predict a UV luminosity density in approximate agreement with both FIREboxHR and observational data at $z\sim{}6-9$, most of them increasingly underestimate the UV luminosity density towards higher redshift. Exceptions are the SPHINX20 simulation (Rosdahl et al., 2022), which matches the predictions of FIREboxHR at $z=10$ but does not provide estimates for higher redshifts, the FIRE-2 zoom-in suites by Ma et al. (2019) and Sun et al. (2023c), which match FIREboxHR out to $z=10$ and $z=12$, respectively, and the FirstLight suite (Ceverino et al., 2024), which underpredicts the observed UV luminosity density at $z\sim{}10-11$ with respect to Donnan et al. (2023a); Donnan et al. (2023b); Donnan et al. (2024) and Finkelstein et al. (2024), but is in approximate agreement with observations at $z\sim{}12-13$.